The following data relates to the market value of economic transactions for the three main sectors of our country's economy.

Sector Value of output Purchases from other Firms
Agriculture 300 160
Manufacturing 200 150
Services 150 120
i. Compute the gross domestic producdt (GDP)
ii. Compute Gross national (GNP) at market prices
iii. Compute Net national Product (NNP) at Market Prices.
iv. Compute Net National Product (NNP) at factor cost.
v. Net domestic product (NDP) at factor cost.
1
Expert's answer
2020-06-09T17:21:16-0400

i) "300+160+200+150+150+120=1080"

ii) "300+200+150=650"

iii)"NNP=GNP+Net\ Tax= 650+0=650"

iv)"-160-150-120=-430"

v) "300-160+200-150+150-120=220"